OpenStudy (anonymous):

on a purchase of $500, how much is saved by taking discounts of 20% and 10%, rather than discounts of 10% and 15%?

OpenStudy (anonymous):

I'm trying to figure out what you mean by taking discounts of 20% and 10% instead of the other two.. How would you get these discounts? after each other.. the differences between the first to.. the discounts of all four?

OpenStudy (anonymous):

the way i was taught was to multiply the first percentage then subtract that number from total number the multiply that number times the second percentage

OpenStudy (anonymous):

multiply the first with what?

OpenStudy (anonymous):

multiply the first percent 20%, for example, with 500

OpenStudy (anonymous):

{80/100 * 90/100 *500} - { 90/100 *85/100 *500}

OpenStudy (anonymous):

I think that depends on how exactly this question is meant.. I can see different solutions.. I'm just not sure what would work in this case.. both methods seem to be possible when the question is asked like this...

OpenStudy (anonymous):

this is the correct one....!!!

OpenStudy (anonymous):

can you multiply the discounts if they are substracted after each other?

OpenStudy (anonymous):

no idon'tknow

OpenStudy (anonymous):

here u need to find the cost after first discount then after that u again have a second discount so u will calculate the cost after second discount but this time u have to calculate the cost after doing discount on the cost which we have got after the first discount

OpenStudy (anonymous):

lets here discuss the first case hope u will understand the second one by urself...!!! so cost after first discount which is of 20% u get 80/100 *500=400 and then we will calculate the price after second discount which is 10% so the cost after second discount will be 90/100*400=360

OpenStudy (anonymous):

hope u understand and tell me what do u get exact answer after u solve....!!

OpenStudy (anonymous):

and lso likewise u have to do in the second case...!! if u dont get just ask me...!!

OpenStudy (anonymous):

thanks i understand you really are a lifesaver

OpenStudy (anonymous):

ur welcome but what u get after solving i wanna make sure u get the correct anwer...!!!

OpenStudy (anonymous):

also. if you want to multiply the total amount with the percentage instead of dividing first.. you have to take the fraction not the percentage... for instance 10% is equal to a fraction of 0,10 and 50% is 0,50

OpenStudy (anonymous):

102.90

OpenStudy (anonymous):

22.50

OpenStudy (anonymous):

that is the same thing you are not subtracting that number from the original answer

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

see u get the price after 2 discounts in the first case=360 lets here discuss the first case hope u will understand the second one by urself...!!! so cost after first discount which is of 20% u get 80/100 *500=400 and then we will calculate the price after second discount which is 10% so the cost after second discount will be 90/100*400=360

OpenStudy (anonymous):

if u get it then i will tell u about the second case

OpenStudy (anonymous):

i have to go now and post another problem. you guys are just awesome and if i could meet you we would be friends forever

OpenStudy (anonymous):

heheeh thanx....!!!!!